OpenStudy (anonymous):

3. The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line? (1 point) y = –x – 5 y = x + 2 y = x – 2 y = –x – 2 4. What is the slope-intercept form of the equation y – 7 = –(x + 4)? (1 point) y – 2 = –(x + 2) y + 7 = –x + y = –+ 2 y = –x – 3 Please help i dont understand it

OpenStudy (anonymous):

basically most of these equations are in the form y=mx+b

OpenStudy (anonymous):

so the m is the slope and the b is the y-intercept

OpenStudy (anonymous):

for 3. you are given b=-2 so the equation must be one of these: y = x – 2 y = –x – 2

OpenStudy (lukebluefive):

The slope is negative: \[\frac{(0- -2)}{-5 - 0} = -\frac{2}{5}\] For your second equation, slope-intercept form looks like this: \[y = mx + b\] So the last option is the only possibility: \[y = -x + 3\]

OpenStudy (anonymous):

now you can plugin the value (-5, 0) because its an xintercept but it can be any other value then calculate the slope by using (y1-y2)/(x1-x2)

OpenStudy (anonymous):

sorry you dont need calculate the slope that way its better to do it like y=mx+b so y=mx-2 now (-5, 0) 0=m(-5)-2 2=-5m m=2/-5

OpenStudy (anonymous):

i dont know why i didnt get any of the options but the equation i got is: y=2/-5 x -2

OpenStudy (anonymous):

sorry @burn23 i cant help you much but i know that this is the way i do it try doing the second one